OpenStudy (anonymous):

Why would the answer when factoring the polynomial 9x^2 − 324y^2 not be (3x-18y)(3x+18y) but instead be 9(x-6y)(x+6y) ?

jimthompson5910 (jim_thompson5910):

because you can factor the GCF 9 from the original polynomial or you can factor two copies of 3 from each group (in parenthesis) to get 3*3 = 9 Either way, the complete factorization will be 9(x-6y)(x+6y)

jimthompson5910 (jim_thompson5910):

by the second method, I mean (3x-18y)(3x+18y) 3(x-6y)3(x+6y) 3*3(x-6y)(x+6y) 9(x-6y)(x+6y)

OpenStudy (anonymous):

how did you get the 6?

jimthompson5910 (jim_thompson5910):

(3x-18y) = 3(x-6y) and (3x+18y) = 3(x+6y)